A soccer player kicks a ball at an angle of 43 ° with a speed of 10m/s. How far from the initial position( position A which the player kicks the ball) will the ball land?
1
Expert's answer
2021-10-15T10:23:38-0400

"L=v^2\sin2\alpha/g=10^2\cdot\sin(2\cdot43°)/9.8=10.18\ (m/s)"